OpenStudy (anonymous):

Tim's phone service charges $25.16 plus an additional $0.21 for each text message sent per month. If Tim's phone bill was $30.41, which equation could be used to find how many text messages, x, Tim sent last month? $0.21x - $25.16 = $30.41 $25.16x - $0.21 = $30.41 $0.21x + $25.16 = $30.41 $25.16x + $0.21 = $30.41

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Tim's phone service charges $25.16 plus an additional $0.21 for each text message sent per month. If Tim's phone bill was $30.41, which equation could be used to find how many text messages, x, Tim sent last month? \(\large\color{slate}{ 25.16 +0.21x= 30.41 }\)

OpenStudy (solomonzelman):

solve this equation for x.. 0.21 is a cost per call, x is the [unknown] number of calls.

OpenStudy (solomonzelman):

first subtract 25.16 from both sides, and then divide both sides by 0.21

OpenStudy (anonymous):

Well I am so dumb that I still don't get it.. Sorry though... :(

OpenStudy (solomonzelman):

\(\large\color{slate}{ 25.16+0.21x=30.41 }\) \(\large\color{slate}{ 25.16+0.21x\color{red}{-25.16}=30.41\color{red}{-25.16} }\) \(\large\color{slate}{ 0.21x=5.25 }\)

OpenStudy (solomonzelman):

can you go from here ?

OpenStudy (anonymous):

25

OpenStudy (solomonzelman):

Yes, there you go!